OpenStudy (anonymous):

\[\sum_{n=1}^{\infty} (-1)^{n} \frac{2^{n}n!}{5*8*11*...*(3n+2)}\] \[\lim_{n \rightarrow \infty} \frac{5*8*11*...*(3n+2)}{2^{n}n!} \frac{2^{n+1}(n+1)!}{5*8*11*...*(3(n+1)+2)}\] \[\lim_{n \rightarrow \infty} \frac{(3n+2)}{2^{n}} \frac{2^{n+1}(n+1)}{(3(n+1)+2)}\] \[\lim_{n \rightarrow \infty} \frac{2^{n+1}}{2^{n}} \frac{(n+1)(3n+2)}{(3(n+1)+2)}\] \[\lim_{n \rightarrow \infty} 2\frac{(n+1)(3n+2)}{(3(n+1)+2)}\]

OpenStudy (anonymous):

hope im on the right track

OpenStudy (anonymous):

Do you want to compute the sum or prove convergence, divergence?

OpenStudy (anonymous):

convergence or divergence

OpenStudy (anonymous):

\[\lim_{n \rightarrow \infty} 2\frac{3n^{2}+5n+2}{3n+5}\]

OpenStudy (anonymous):

Limits is infinity

OpenStudy (anonymous):

because the numerator is a degree higher than the denominator?

OpenStudy (anonymous):

I think you can cancel the (3n+2) in the numerator cause in the denominator the sequence should be 5*8*11*...*(3n+2)*(3n+5), don't you?

OpenStudy (anonymous):

I considered that ...but is it legal?

OpenStudy (anonymous):

I think it's legal, that makes the limit 2/3 < 1 and the serie converges

OpenStudy (anonymous):

\[\lim_{n \rightarrow \infty} \frac{(3n+2)}{2^{n}} \frac{2^{n+1}(n+1)}{(3(n+1)+2)}\] show me the math

OpenStudy (anonymous):

please

OpenStudy (anonymous):

because according to eliassaab it's divergent

OpenStudy (anonymous):

Before that on the second step

OpenStudy (anonymous):

or am is misunderstanding @eliassaab

OpenStudy (anonymous):

Ok, maybe I'm wrong then

OpenStudy (anonymous):

Your series is convergent. Try to see why.

OpenStudy (anonymous):

why?

OpenStudy (anonymous):

lolz sorry guys...i'm not quite understanding

OpenStudy (anonymous):

\[\lim_{n \rightarrow \infty} 2\frac{3n^{2}+5n+2}{3n+5}\] am i right up to this point?

OpenStudy (anonymous):

I think you're not, I mean on second step, we have in the numerator 5*8*11*...*(3n+2) right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Ok, in the denominator we have the x sub(n+1) and it is 5*8*11*...*(3n+2)*(3n+5)

OpenStudy (anonymous):

\[ \lim_{n->\infty } \left |\frac {a_{n+1}}{a_n} \right|=\frac 23 \]

OpenStudy (anonymous):

We can cancel the (3n+2) terms, that what I'm trying to say, but as I said above, maybe I'm wrong

OpenStudy (anonymous):

So it is convergent.

OpenStudy (anonymous):

\[\lim_{n \rightarrow \infty} 2\frac{(n+1)(3n+2)}{3(n+1)+2}\] @eliassaab does it matter that the numerator is a degree higher than the denominator?

OpenStudy (anonymous):

ok you are both right according to my solution manual...

OpenStudy (anonymous):

but i'm still having trouble with the elimination part I guess

OpenStudy (anonymous):

sigh...the mystery \[\lim_{n->\infty } \left |\frac {a_{n+1}}{a_n} \right | = \lim_{n->\infty } \frac {a_{n+1}}{a_n} =\\ \lim_{n->\infty } \frac {2^{n+1} (n+1)!} {5. 8 \cdots (3n+1)(3n+4)} \frac{5. 8 \cdots (3n+1)} {2^{n} n!}\\\] what happened here \[ =\frac {2 n}{3n+4}\to \frac 2 3\]

OpenStudy (anonymous):

Did you get it?

OpenStudy (anonymous):

i get this part \[=\frac {2 n}{3n+4}\to \frac 2 3\]

OpenStudy (anonymous):

By the ration test, the series converges.

OpenStudy (anonymous):

@knock @eliassaab plz don't hate me lolz ...i'm not intentionally trying to be so stubborn...i guess i'll just stare at it a little longer

OpenStudy (anonymous):

is it 2/3 or -2/3

OpenStudy (anonymous):

You have to take the absolute value of the ratio. So it is 2/3

OpenStudy (anonymous):

i understand it's convergent but i guess i'm having trouble with the algebra part eliminating form the numerator and denominator...(which is the easiest part of this problem)

OpenStudy (anonymous):

What is (n+1)!/n!

OpenStudy (anonymous):

n+1

OpenStudy (anonymous):

Imagine that you have this sequence 5*8*11...(3n+2), 3n+2 verify the terms because for n=1 it's 5, for n=2 it's 8 and so well then you see the diference between each number it's 3, so the next term after 3n+2 should be 3n+5, as all terms are multipliying to each other we can cancel all of the and we have 3n+5 in the denominator

OpenStudy (anonymous):

I had a small misprint.

OpenStudy (anonymous):

ok now we can cross out 3n+1

OpenStudy (anonymous):

VICTORY!!!!

OpenStudy (anonymous):

We can cross out 5.8, ... (3n+1)

OpenStudy (anonymous):

yep yep now I can see it

OpenStudy (anonymous):

Great ^^

OpenStudy (anonymous):

Great

OpenStudy (anonymous):

thanks! saved my life you two

OpenStudy (anonymous):

Yep but I think the terms are 3n+2 aren't they? Cause 3n+4 doesn't makes 5*8*11... for n=1,2,...

OpenStudy (anonymous):

It is 3n +5

OpenStudy (anonymous):

Ok, agree then yw guys

OpenStudy (anonymous):

\[ \lim_{n->\infty } \left |\frac {a_{n+1}}{a_n} \right | = \lim_{n->\infty } \frac {a_{n+1}}{a_n} =\\ \lim_{n->\infty } \frac {2^{n+1} (n+1)!} {5. 8 \cdots (3n+2)(3n+5)} \frac{5. 8 \cdots (3n+2)} {2^{n} n!}\\ =\frac {2( n+1)}{3n+5}\to \frac 2 3 \]

OpenStudy (anonymous):

makes sense now

OpenStudy (anonymous):

The numerical sums of your series is -0.269723